OpenStudy (anonymous):

Solve V = π r2h for h

Parth (parthkohli):

Divide both sides by \(\pi r^2\)

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

that's it, you divide both sides by pi r^2 and you get yoru answer

OpenStudy (anonymous):

You also have to subtract the v from both sides. h = pir^2-V

OpenStudy (anonymous):

@mialarson03 ParthKohli and PaulK are correct. You don't have to subtract V from both sides because if you do so, Pi (r^2) h - V = 0 Now if you divide both sides by Pi (r^2), you will get h - V/pi (r^2) = 0 or h = V/pi (r^2) You go a "big" circle and come back to the same point. All you have to do is, as the other two said, divide both sides by pi (r^2), you get V/pi (r^2) = pi (r^2) h / pi (r^2) = h

OpenStudy (anonymous):

thnk you all